# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems

### v2

@@ Abstract.
In this work, we study the applicability of the Azouani-Olson-Titi (AOT) nudging algorithm for continuous data assimilation to evolutionary dynamical systems that are not dissipative. Specifically, we apply the AOT algorithm a partially dissipative variant of the Lorenz 1963 system, the Korteweg-de Vries (KdV) in 1D, and the 2D incompressible Euler equations. Our analysis reveals that both the Euler and KdV equations lack the finitely many determining modes property, leading to the construction of infinitely many solutions with exactly the same sparse observational data, which data assimilation methods cannot distinguish between. Simultaneously, we numerically verify that the AOT algorithm successfully recovers these counterexamples for the damped and driven KdV equation, which is dissipative. Additionally, to further support our argument, we present numerical evidence showing that the AOT algorithm is ineffective at accurately recovering solutions for a partially dissipative variant of the Lorenz 1963 system.
@@ 1. Introduction
Accurate computational prediction of nonlinear evolution systems encounters two major challenges. That of initialization and of long-time accuracy, namely, how do we initialize the system and how do we ensure that our computed solution remains accurate for long intervals of time. Data assimilation is a class of techniques that attempts to resolve both of these issues by incorporating observational data into the underlying evolutionary physical model. The purpose of data assimilation is to alleviate the reliance on knowing the complete state of the initial data and achieve synchronization of our simulated solution with the true dynamics by utilizing adequately collected sparse spatio-temporal observational data (measurements).
@@
In this work, we are concerned with the application of data assimilation algorithms to systems that are not known to be dissipative. By dissipative, we mean that equations possess a compact absorbing ball or set in a relevant vector space (e.g. $L^{2}$ or $H^{1}$), which is a set that all solutions to the system eventually enter into and remain within for all future times. The existence of such an absorbing ball (or set) can be used to show the existence of a global attractor with finite fractal dimension, which in turns yields a finite number of parameters that determine the asymptotic in time behavior of solutions for dissipative dynamical systems. One relevant example of data assimilation being applied to non-dissipative dynamical systems is that of the shallow water equations, which have long been used as a simplified model for the ocean dynamics. Contrary to the case of dissipative dynamical systems [23], it is unknown whether these equations have finitely many determining parameters, such as finitely many Fourier modes or nodal values which determine the solution uniquely. Yet researchers still apply data assimilation to these systems, see, e.g., [13, 56, 77]. The property, for example, of finitely many determining Fourier modes [37, 54, 23] is an essential first litmus test for data assimilation to work, as it ensures the existence of a minimum length scale at which measurements and observations determine the asymptotic behavior of the system we wish to predict. Without this property, there is no guarantee that any amount of observations will be sufficient to recover the dynamics of the true solution one wishes to predict.
@@
This work focuses on characterizing the class of systems for which one should not expect the implementation of the Azouani-Olson-Titi (AOT) continuous data assimilation algorithm, introduced in [4, 5], is able to effectively recover. Since its development, this algorithm has been the subject of much study, both analytically [1, 6, 7, 8, 9, 10, 16, 18, 21, 20, 26, 28, 31, 32, 33, 34, 35, 38, 39, 42, 41, 43, 50, 51, 52, 53, 58, 65, 66, 68, 70, 73, 81] and computationally [2, 19, 22, 25, 40, 44, 59, 60, 63, 57]. We note that the AOT algorithm has been referenced a few times in the literature under the names interpolated nudging or simply nudging, among others. In the context of classical methods of data assimilation, nudging refers to a method developed by Anthes and Hokes [3, 49] in the 1970’s for numerical weather prediction. The AOT algorithm is superficially similar to nudging, but is technically and conceptually different. In the classical algorithm of Anthes and Hokes one nudges the system only at the discrete nodal points of the observational data. On the other hand, the AOT algorithm works by constructing a spatial approximation interpolation operator based on the observational data and then nudges at every point in the vicinity of the observation nodal point using the interpolation operator. Nevertheless, for simplicity, we will refer to the AOT algorithm as “nudging” for the remainder of this paper. Notably, the AOT data assimilation algorithm was designed for dissipative evolutionary nonlinear systems motivated by, and capitalizing on, the fact that the long-term dynamics of such systems are determined by the dynamics of large spatial scales (see, e.g., [23] and the references therein). Such systems include for example the 2D Navier-Stokes equations, the 2D Rayleigh-Bernard convection problem, as well as the 1D Kuramoto-Sivashinsky equation, and some nonlinear reaction-diffusion systems, to name some (cf. [76]).
@@
We note that the notion of dissipation for dynamical systems is often conflated with the inclusion of diffusion and viscous effects. We note that inclusion of a Laplacian term is one way to obtain a dissipative system as we have described, however dissipation can also be realized through damping terms, as in the Korteweg-de Vries equation (see e.g. [46, 45]). We note that for many dynamical systems relevant to geophysics, the model equations are assumed to be not be dissipative. In practice, modelers often re-introduce small artificial diffusion terms purely for numerical stability. We note however that the inclusion of such terms may guarantee that the model equations may be dissipative even though the underlying original system is not.
@@
In this work we consider a few illustrative examples of non-dissipative dynamical systems in order to demonstrate that AOT data assimilation algorithm cannot be utilized to accurately recover the true solution. We want to illustrate that in the case of the PDEs we consider here, that the failure of the AOT algorithm, but rather is an intrinsic property of the equations considered here due to the convolutions structure of the nonlinearity. By this we mean that the pathologic solutions we consider here present an issue for any method of data assimilation, as the solutions are not uniquely identifiable from a fixed number of observed Fourier modes.
@@
In this work, we will consider three examples of non-dissipative dynamical systems; a partially dissipative variant of the 1963 Lorenz system, the classical Korteweg-de Vries equation in 1D, and the incompressible Euler equations in 2D. The 1963 Lorenz system is a chaotic system of coupled ODEs that has been used to test methods of data assimilation. We note that the AOT algorithm has applied to the fully dissipative 1963 Lorenz system in [11, 71, 17]. We consider the KdV equation in particular, as it was proven in [45, 46] that the system has a finite-dimensional global attractor in the presence of a damping and forcing terms. Moreover, it was shown in [53] that the nudging algorithm recovers the true solution given sufficiently many observed Fourier modes. Without the presence of forcing and damping, the KdV equation is a dispersive equation that has infinitely many conserved quantities, the $L^{2}$ norm being among them. The lack of dissipation and these conservation properties make this equation ideal to consider as a relatively simple 1D model to explore why data assimilation fails to recover solutions to systems without finitely many determining modes. We note that while we do consider the KdV equation, the analytical findings present here extend to a broader class of PDEs that conserve the $L^{2}$ norm featuring the same nonlinearity.
@@
The remainder of this manuscript is organized as follows: in Section˜2 we consider a partially dissipative variant of the Lorenz system and we show that nudging fails to recover the reference solution in this setting. We begin our discussion of the application of nudging to recover the solutions to the KdV equation in Section˜3.1, where we discuss various mathematical preliminaries regarding the formulation of the KdV equation and we show that, unlike the damped and driven case, the standard version of KdV does not have the finitely many determining modes property. In Section˜3.2 we show the numerical construction of pathological examples where nudging fails to recover the true solution for the KdV equations and we additionally verify the validity of the AOT nudging algorithm for these counterexamples when the system is dissipative. In Section˜4 we extend our results from the KdV regime to the case of 2D incompressible Euler equations. In Section˜4.1 we outline the analytical construction of pathological counterexamples for this regime and in Section˜4.2 we demonstrate these findings computationally. We end this work in Section˜5 with a discussion of the conclusions we reach from our numerical results.
@@ 2. Lorenz 1963 System
As the focus of this work is examining the applicability of the AOT algorithm to dynamical systems that are not dissipative, we now turn to examine a nondissipative variant of the Lorenz 1963 system. Lorenz introduced in 1963 a simplified/reduced model of atmospheric convection [62], and so it is of particular interest for studying methods of data assimilation, see e.g. [48, 11, 28, 47, 69, 61] and the references within. The 1963 Lorenz system is given by:
@@
It is also known that applying the AOT algorithm, or a synchronization scheme, e.g., [48, 47], by observing only the $X$ component is sufficient to synchronize the nudged solution with the reference solution (see e.g. [48, 11, 47]). In particular, we can consider the following three different synchronization (nudging) schemes from [47] (which was motivated by [67]) to understand when data assimilation works for this system.
@@
In the above systems, Equations˜2, 3 and 4, correspond to the regimes when the observed values of $X$, $Y$, or $Z$, respectively, are observed and inserted directly into the equations for the remaining, unobserved components. That is, e.g. in Equation˜2 we observe the true value of $X$ continuously in time and insert that into the evolution equations of $y$ and $z$. In these equations we note that numerically we should not be simulating the evolution of $X$, $Y$, or $Z$ as they appear in Equations˜2, 3 and 4, respectively, but rather these values are observed (given) from the true solution in Equation˜1 and are inserted directly into the equations for the remaining variables. Analysis on Equations˜2, 3 and 4 was conducted in [47] where it was shown that the continuous coupling on $X$ or $Y$ alone was sufficient to obtain synchronization between the assimilated equations and the reference solution, however the coupling seen in Equation˜4 was not sufficient to obtain convergence in either the $x$ or $y$ components to the corresponding components $X$ or $Y$ of the true solution. We do note that in [47] the equations for Equation˜1 were written in a slightly different form, however this can be easily reconciled by using the transformation $\tilde{Z}=Z+(r+\sigma)$.
For the sake of completeness, we will discuss the convergence results for the above systems, Equations˜2, 3 and 4 proven in [47] and discuss how the results break down in the absence of dissipation. Let
$\delta=(X-x,Y-y,Z-z),$ (5)
where $(X,Y,Z)$ is the given solution of the original Lorenz system Equation˜1 and $(x,y,z)$ is the solution to one of the assimilated equation, Equations˜2, 3 and 4. Here we define $\delta$ with the understanding that the variable in $(x,y,z)$ missing from the assimilated equation is taken to be equal to the corresponding variable in Equation˜1, e.g. when considering Equation˜2 we define $x(t)\mathrel{\mathop{\ordinarycolon}}=X(t)$ for all time, $t$ and have $\delta=(0,Y-y,Z-z)$. The convergence results for the synchronization algorithms, Equations˜2, 3 and 4 are given as follows:
@@
$\mathinner{\!\left\lVert\delta(t)\right\rVert}^{2}_{\ell^{2}}\leq\mathinner{\!\left\lVert\delta_{0}\right\rVert}_{\ell^{2}}e^{-2t}.$ (6)
@@
$\begin{split}\mathinner{\!\left\lVert\delta(t)\right\rVert}^{2}_{\ell^{2}}&\leq|X(0)-x(0)|e^{-2\sigma}+|Z(0)-z(0)|e^{-bt}\\ &\quad+\frac{J|X(0)-x(0)|^{2}}{b(2\sigma-b)}\left(e^{-bt}-e^{-2\sigma t}\right),\end{split}$ (7)
@@
We note that the analysis done in the proof of Theorem˜2.1 is dependent on the particular parameters and requires $\sigma>0$, $b>0$ and $r>0$. To illustrate how these sorts of estimates break down in the absence of dissipation, we consider the case when both $X$ and $Y$ are observed, i.e., are given from the true solution of Equation˜1. This leads to the following system of equations
@@
We note that the above discussion has been on the synchronization data assimilation scheme, given by the direct insertion, as it was first introduced in [67], of the observed quantities into the model equations. We now consider the nudged Lorenz system for the observed $X$ and $Y$ components of Equation˜1, given as follows:
@@
At any fixed time, $t$, the solutions, $(x(t;\mu),y(t;\mu),z(t;\mu))$, to the nudged system, Equation˜13, are expected to converge to the solution $(X(t),Y(t),z(t))$ of the synchronization algorithm, given by Equation˜10 as $\mu\to\infty$. This can be justified rigorously by applying the argument present in [15] to the Lorenz 1963 system, provided both methods are initialized with the same initial data. An analogous statement was made in [71] for the Lorenz system in particular (with $b>0$), where it was shown that the solution, $x(t;\mu)$, of Equation˜13, converges to $X(t)$ at any fixed time, $t$, as we send $\mu\to\infty$ when the nudging term is present only on the equation for $x$ in Equation˜13. In the case where $b=0$, if we assume a priori that the solutions of Equation˜13 converge to those of Equation˜10 as $\mu\to\infty$, one can deduce using the reverse triangle inequality, that $\mathinner{\!\left\lVert\delta(t)\right\rVert}^{2}_{\ell^{2}}\geq|z_{0}-Z_{0}|^{2}$, where $\delta$ is given as defined previously in Equation˜5. Therefore, we expect the nudged Lorenz system, Equation˜13, to fail to recover the true solution in the same way that Equation˜10 fails when dissipation is removed from the equations for $z$.
@@
We have numerically implemented the Lorenz system as in [11, 71] using an explicit Euler scheme for the time discretization. As in [71], we will utilize $\Delta t=1e-4$ for our timestep. We note that in many works the reference solution is initialized by running the solution forward in time until it is approximately on the global attractor (see e.g. [11, 48]), however as we are attempting to test nudging for non-dissipative or partially dissipative systems, which need not possess global attractors, we will instead be utilizing the initial data $(X,Y,Z)=(30,40,50)$ as in [71].
First, we verify the validity of nudging on the Lorenz system with full dissipation. Using the initial data and parameter choices discussed previously we apply nudging to the Lorenz system with $b=8/3$. The results can be seen in Figure˜1, where we have applied nudging to only the first two components of Equation˜13. Here we can see that all variables converge to an error level of approximately 1e121e-12 with spurious dips below this error level in each component. Note that here we see exponential convergence to the true solution, as we expect in this case.
@@
Next, we examine what happens when we apply nudging for the partially dissipative Lorenz system, i.e., with $b=0$. The results can be seen in Figure˜2. Note that here we still see exponential decay to the true solution in both $x$ and $y$. While nudging appears to affect $z$ initially, this effect quickly fades after a short interval of time and the error appears to stabilize to be approximately constant $1e+1$. The error level that is attained appears to be directly linked to the difference in the initial conditions for the $z$ component. To illustrate this, we plotted the $\ell^{2}$ (Euclidean norm) error over time when both the reference and nudged equations are initialized with the same initial data, except for the $z$ component. Here $Z(t_{0})=50$ and $z(t_{0})=50+\delta$, where $\delta$ ranges in values from 1e11e-1 to 1e121e-12. The results of these tests can be seen in Figure˜3 and they illustrate that the $\ell^{2}$ error is highly sensitive to small perturbations in the initial condition for $z$. We note that the behavior seen in Figure˜3 appears to align with what one should expect from our earlier discussion of Equation˜11 for the synchronization algorithm. We note that we obtain an initial exponential decay in the error in the $z$ component which is driven by the convergence in both $x$ and $y$. This initial convergence levels off and becomes approximately constant over time.
@@
Figure 3. Effect of perturbations of the true initial condition on convergence for nudging for Lorentz. Here $Z(t_{0})=50$ and $z(t_{0})=50+\delta$, with $X(t_{0})=x(t_{0})=20$ and $Y(t_{0})=y(t_{0})=30$ with $\mu=10$. Parameters used were $\sigma=10$, $\rho=28$, and $b=0$.

### v3

@@ Abstract.
In this work, we study the applicability of the Azouani-Olson-Titi (AOT) nudging algorithm for continuous data assimilation to evolutionary dynamical systems that are not dissipative. Specifically, we apply the AOT algorithm to a partially dissipative variant of the Lorenz 1963 system, the Korteweg-de Vries equation (KdV) in 1D, and the 2D incompressible Euler equations. Our analysis reveals that both the Euler and KdV equations lack the finitely many determining modes property, leading to the construction of infinitely many solutions with exactly the same sparse observational data, which data assimilation methods cannot distinguish between. Simultaneously, we numerically verify that the AOT algorithm successfully recovers these counterexamples for the damped and driven KdV equation, which is dissipative. Additionally, to further support our argument, we present numerical evidence showing that the AOT algorithm is ineffective at accurately recovering solutions for a partially dissipative variant of the Lorenz 1963 system.
@@ 1. Introduction
Accurate computational prediction of nonlinear evolution systems encounters two major challenges: initialization and long-time accuracy. These refer to the difficulty in accurately initializing the system and ensuring that our computed solution remains accurate for long intervals of time, respectively. Data assimilation is a class of techniques that attempts to resolve both of these issues by incorporating observational data into the underlying evolutionary physical model. The purpose of data assimilation is to alleviate the reliance on knowing the complete state of the initial data and achieve synchronization of our simulated solution with the true dynamics by using adequately collected sparse spatio-temporal observational data (measurements).
@@
In this work, we are concerned with the application of data assimilation algorithms to systems that are not known to be dissipative. By dissipative, we mean that equations possess a compact absorbing ball or set in a relevant vector space (e.g. $L^{2}$ or $H^{1}$), which is a set that all solutions to the system eventually enter into and remain within for all future times. The existence of such an absorbing ball (or set) can be used to show the existence of a global attractor with finite fractal dimension, which in turn yields a finite number of parameters that determine the asymptotic in time behavior of solutions for dissipative dynamical systems. One relevant example of data assimilation being applied to non-dissipative dynamical systems is that of the shallow water equations, which have long been used as a simplified model for the ocean dynamics. Contrary to the case of dissipative dynamical systems [23], it is unknown whether these equations have finitely many determining parameters, such as finitely many Fourier modes or nodal values which determine the solution uniquely. Yet researchers still apply data assimilation to these systems, see, e.g., [13, 56, 77]. The property, for example, of finitely many determining Fourier modes [37, 54, 23] is an essential first litmus test for data assimilation to work, as it ensures the existence of a minimum length scale at which measurements and observations determine the asymptotic behavior of the system we wish to predict. Without this property, there is no guarantee that any amount of observations will be sufficient to recover the dynamics of the true solution one wishes to predict.
@@
This work focuses on characterizing the class of systems for which one should not expect the implementation of the Azouani-Olson-Titi (AOT) continuous data assimilation algorithm, introduced in [4, 5], to be effective. Since its development, this algorithm has been the subject of much study, both analytically [1, 6, 7, 8, 9, 10, 16, 18, 21, 20, 26, 28, 31, 32, 33, 34, 35, 38, 39, 42, 41, 43, 50, 51, 52, 53, 58, 65, 66, 68, 70, 73, 81] and computationally [2, 19, 22, 25, 40, 44, 59, 60, 63, 57]. The AOT algorithm has been referenced a few times in the literature under the names interpolated nudging or simply nudging, among others. In the context of classical methods of data assimilation, nudging refers to a method developed by Anthes and Hoke [3, 49] in the 1970’s for numerical weather prediction. The AOT algorithm is superficially similar to nudging, but is technically and conceptually different. In the classical algorithm of Anthes and Hoke one nudges the system only at the discrete nodal points of the observational data. On the other hand, the AOT algorithm works by constructing a spatial approximation interpolation operator based on the observational data and then nudges at every point in the vicinity of the observation nodal point using the interpolation operator. Nevertheless, for simplicity, we will refer to the AOT algorithm as “nudging” for the remainder of this paper. Notably, the AOT data assimilation algorithm was designed for dissipative evolutionary nonlinear systems motivated by, and capitalizing on, the fact that the long-term dynamics of such systems are determined by the dynamics of large spatial scales (see, e.g., [23] and the references therein). Such systems include, for example, the 2D Navier-Stokes equations, the 2D Rayleigh-Bénard convection problem, as well as the 1D Kuramoto-Sivashinsky equation, and some nonlinear reaction-diffusion systems, to name some (cf. [76]).
@@
The notion of dissipation for dynamical systems is often conflated with the inclusion of diffusion and viscous effects. Including a Laplacian term is one way to obtain a dissipative system; however, dissipation can also be realized through damping terms, as in the Korteweg-de Vries equation (see e.g. [46, 45]). For many dynamical systems relevant to geophysics, the model equations are assumed to be non-dissipative. In practice, modelers often re-introduce small artificial diffusion terms purely for numerical stability. We note however that the inclusion of such terms may guarantee that the model equations may be dissipative even though the underlying original system is not.
@@
In this work we consider a few illustrative examples of non-dissipative dynamical systems in order to demonstrate that the AOT data assimilation algorithm cannot be used to accurately recover the true solution. We emphasize that our results are not isolated specifically to the AOT algorithm, but rather extend generically to data assimilation algorithms in general. The failure of data assimilation algorithms to recover solutions to non-dissipative dynamical systems arises as an intrinsic property of the particular equations considered. By this we mean that the pathological solutions we consider here present an issue for any method of data assimilation, as the solutions are not uniquely identifiable from a fixed number of observed Fourier modes.
@@
In this work, we consider three examples of non-dissipative dynamical systems; a partially dissipative variant of the 1963 Lorenz system, the classical Korteweg-de Vries equation in 1D, and the incompressible Euler equations in 2D. The 1963 Lorenz system is a chaotic system of coupled ODEs that has been used to test methods of data assimilation, including the AOT algorithm in the fully dissipative regime (see e.g. [11, 71, 17]). For the 1963 Lorenz system we show that removing the dissipation from even a single component is sufficient to prevent the AOT algorithm from recovering the true solution. We consider the KdV equation in this work as, in particular, as it was proven in [45, 46] that the system has a finite-dimensional global attractor in the presence of a damping and forcing terms. Moreover, it was shown in [53] that the nudging algorithm recovers the true solution given sufficiently many observed Fourier modes. Without the presence of forcing and damping, the KdV equation is a dispersive equation that has infinitely many conserved quantities, the $L^{2}$ norm being among them. The lack of dissipation and these conservation properties make this equation ideal to consider as a relatively simple 1D model to explore why data assimilation fails to recover solutions to systems without finitely many determining modes. While we do consider the KdV equation, the analytical findings present here extend to a broader class of PDEs that conserve the $L^{2}$ norm featuring the same nonlinearity. To illustrate this, we also consider the incompressible Euler equations in 2D, which also conserve the $L^{2}$ norm and are not known to have finitely many determining modes. The 2D incompressible Euler equations are of particular interest as they are the inviscid counterpart to the 2D incompressible Navier-Stokes equations, which are a classical example of a dissipative system and was the original setting of the AOT algorithm [4].
@@
The remainder of this manuscript is organized as follows: in Section˜2 we consider a partially dissipative variant of the Lorenz system and show that nudging fails to recover the reference solution in this setting. We begin our discussion of the application of nudging to recover the solutions to the KdV equation in Section˜3.1, where we discuss various mathematical preliminaries regarding the formulation of the KdV equation and show that, unlike the damped and driven case, the standard version of KdV does not have the finitely many determining modes property. In Section˜3.2 we show the numerical construction of pathological examples where nudging fails to recover the true solution for the KdV equations and we additionally verify the validity of the AOT nudging algorithm for these counterexamples when the system is dissipative. In Section˜4 we extend our results from the KdV regime to the case of 2D incompressible Euler equations. In Section˜4.1 we outline the analytical construction of pathological counterexamples for this regime and in Section˜4.2 we demonstrate these findings computationally. We end this work in Section˜5 with a discussion of the conclusions we reach from our numerical results.
@@ 2. Lorenz 1963 System
As the focus of this work is examining the applicability of the AOT algorithm to dynamical systems that are not dissipative, we now examine a nondissipative variant of the Lorenz 1963 system. Lorenz introduced in 1963 a simplified/reduced model of atmospheric convection [62], and so it is of particular interest for studying methods of data assimilation, see e.g. [48, 11, 28, 47, 69, 61] and the references within. The 1963 Lorenz system is given by:
@@
It is also known that applying the AOT algorithm, or a synchronization scheme, (see e.g., [48, 47]), by observing only the $X$ component is sufficient to synchronize the nudged solution with the reference solution (see e.g. [48, 11, 47]). In particular, we can consider the following three different synchronization (nudging) schemes from [47] (which was motivated by [67]) to understand when data assimilation works for this system.
@@
In the above systems, Equations˜2, 3 and 4, correspond to the regimes when the observed values of $X$, $Y$, or $Z$ are inserted directly into the equations for the remaining, unobserved components. That is, e.g. in Equation˜2 we observe the true value of $X$ continuously in time and insert that into the evolution equations of $y$ and $z$. In these equations we note that numerically we should not be simulating the evolution of $X$, $Y$, or $Z$ as they appear in Equations˜2, 3 and 4, respectively, but rather these values are observed (given) from the true solution in Equation˜1 and are inserted directly into the equations for the remaining variables. Analysis on Equations˜2, 3 and 4 was conducted in [47], where it was shown that the continuous coupling on $X$ or $Y$ alone was sufficient to obtain synchronization between the assimilated equations and the reference solution; however, the coupling seen in Equation˜4 was not sufficient to obtain convergence in either the $x$ or $y$ components to the corresponding components $X$ or $Y$ of the true solution. We do note that in [47] the equations for Equation˜1 were written in a slightly different form, however this can be easily reconciled by using the transformation $\tilde{Z}=Z+(r+\sigma)$.
For the sake of completeness, we will discuss the convergence results for the above systems, Equations˜2, 3 and 4 proven in [47] and how the results break down in the absence of dissipation. Let
w=(Xx,Yy,Zz),w=(X-x,Y-y,Z-z), (5)
where $(X,Y,Z)$ is the given solution of the original Lorenz system Equation˜1 and $(x,y,z)$ is the solution to one of the assimilated equation, Equations˜2, 3 and 4. Here we define $w$ with the understanding that the variable in $(x,y,z)$ missing from the assimilated equation is taken to be equal to the corresponding variable in Equation˜1, e.g. when considering Equation˜2 we define $x(t)\mathrel{\mathop{\ordinarycolon}}=X(t)$ for all time, $t$ and have w=(0,Yy,Zz)w=(0,Y-y,Z-z). The convergence results for the synchronization algorithms, Equations˜2, 3 and 4 are given as follows:
@@
$\mathinner{\!\left\lVert w(t)\right\rVert}^{2}_{\ell^{2}}\leq\mathinner{\!\left\lVert w_{0}\right\rVert}_{\ell^{2}}e^{-2t}.$ (6)
@@
$\begin{split}\mathinner{\!\left\lVert w(t)\right\rVert}^{2}_{\ell^{2}}&\leq|X(0)-x(0)|e^{-2\sigma}+|Z(0)-z(0)|e^{-bt}\\ &\quad+\frac{J|X(0)-x(0)|^{2}}{b(2\sigma-b)}\left(e^{-bt}-e^{-2\sigma t}\right),\end{split}$ (7)
@@
The analysis done in the proof of Theorem˜2.1 is dependent on the particular parameters and requires $\sigma>0$, $b>0$, and $r>0$. To illustrate how these sorts of estimates break down in the absence of dissipation, we consider the case when both $X$ and $Y$ are observed, i.e., are given from the true solution of Equation˜1. This leads to the following system of equations
@@
The above discussion has been on the synchronization data assimilation scheme, given by the direct insertion, as it was first introduced in [67], of the observed quantities into the model equations. We now consider the nudged Lorenz system for the observed $X$ and $Y$ components of Equation˜1, given as follows:
@@
At any fixed time, $t$, the solutions, $(x(t;\mu),y(t;\mu),z(t;\mu))$, to the nudged system, Equation˜13, are expected to converge to the solution $(X(t),Y(t),z(t))$ of the synchronization algorithm, given by Equation˜10 as $\mu\to\infty$. This can be justified rigorously by applying the argument in [15] to the Lorenz 1963 system, provided both methods are initialized with the same initial data. An analogous statement was made in [71] for the Lorenz system in particular (with $b>0$), where it was shown that the solution, $x(t;\mu)$, of Equation˜13, converges to $X(t)$ at any fixed time, $t$, as we send $\mu\to\infty$ when the nudging term is present only on the equation for $x$ in Equation˜13. In the case where $b=0$, if we assume a priori that the solutions of Equation˜13 converge to those of Equation˜10 as $\mu\to\infty$, one can deduce using the reverse triangle inequality, that $\mathinner{\!\left\lVert w(t)\right\rVert}^{2}_{\ell^{2}}\geq|z_{0}-Z_{0}|^{2}$, where $w$ is given as defined previously in Equation˜5. Therefore, we expect the nudged Lorenz system, Equation˜13, to fail to recover the true solution in the same way that Equation˜10 fails when dissipation is removed from the equations for $z$.
@@
The Lorenz system is implemented numerically as in [11, 71] using an explicit Euler scheme for the time discretization. All simulations depicted in this study were performed Matlab (R2023a) using our own code. As in [71], we will utilize $\Delta t=10^{-4}$ for our timestep. In many works the reference solution is initialized by running the solution forward in time until it is approximately on the global attractor (see e.g. [11, 48]), however as we are attempting to test nudging for non-dissipative or partially dissipative systems, which need not possess global attractors, we will instead be using the initial data $(X,Y,Z)=(30,40,50)$ as in [71].
First, we verify the validity of nudging on the Lorenz system with full dissipation. Using the initial data and parameter choices discussed previously we apply nudging to the Lorenz system with $b=8/3$. The results can be seen in Figure˜1, where we have applied nudging to only the first two components of Equation˜13. We see that all variables converge to an error level of approximately $10^{-12}$ with dips below this error level in each component. These dips are expected due to the oscillations of the reference solution. Note that here we see exponential convergence to the true solution, as we expect in this case.
@@
Next, we examine what happens when we apply nudging for the partially dissipative Lorenz system, i.e., with $b=0$. The results can be seen in Figure˜2. Note that here we still see exponential decay to the true solution in both $x$ and $y$. While nudging appears to affect $z$ initially, this effect quickly fades after a short interval of time and the error appears to stabilize to be approximately constant $10^{1}$. The error level that is attained appears to be directly linked to the difference in the initial conditions for the $z$ component. To illustrate this, we plotted the $\ell^{2}$ (Euclidean norm) error over time when both the reference and nudged equations are initialized with the same initial data, except for the $z$ component. Here $Z(t_{0})=50$ and $z(t_{0})=50+\delta$, where $\delta$ ranges in values from $10^{-1}$ to $10^{-12}$. The results of these tests can be seen in Figure˜3 and they illustrate that the $\ell^{2}$ error is highly sensitive to small perturbations in the initial condition for $z$. The behavior seen in Figure˜3 appears to align with what one should expect from our earlier discussion of Equation˜11 for the synchronization algorithm. We obtain an initial exponential decay in the error in the $z$ component which is driven by the convergence in both $x$ and $y$. This initial convergence levels off and becomes approximately constant over time.
@@
Figure 3. Effect of perturbations of the true initial condition on convergence for nudging for Lorenz. Here $Z(t_{0})=50$ and $z(t_{0})=50+\delta$, with $X(t_{0})=x(t_{0})=20$ and $Y(t_{0})=y(t_{0})=30$ with $\mu=10$. Parameters used were $\sigma=10$, $\rho=28$, and $b=0$.